# VisualRWKV: Exploring Recurrent Neural Networks for Visual Language Models

### v1

@@ Abstract
Visual Language Models (VLMs) have rapidly progressed with the recent success of large language models. However, there have been few attempts to incorporate efficient linear Recurrent Neural Networks (RNNs) architectures into VLMs. In this study, we introduce VisualRWKV, the first application of a linear RNN model to multimodal learning tasks, leveraging the pre-trained RWKV language model. We propose a data-dependent recurrence and sandwich prompts to enhance our modeling capabilities, along with a 2D image scanning mechanism to enrich the processing of visual sequences. Extensive experiments demonstrate that VisualRWKV achieves competitive performance compared to Transformer-based models like LLaVA-1.5 on various benchmarks. To facilitate further research and analysis, we have made the checkpoints and the associated code publicly accessible at the following GitHub repository: https://github.com/howard-hou/VisualRWKV.
@@
Haowen Hou  and Peigen Zeng  and Fei Ma  and Fei Richard Yu Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ), Shenzhen, China {houhaowen, yufei}@gml.ac.cn https://github.com/howard-hou/VisualRWKV
@@ 3.1 Preliminaries
As shown in Figure 3, trainable variable $\mu_{g}$, $\mu_{r}$, $\mu_{k}$, $\mu_{v}$ are used in a linear combination of $x_{t}$ and $x_{t-1}$, to achieve a simple time mixing, which interpolate between the inputs of the current and previous time-steps. The combination of shifted previous step and current step was linear projected through projection matrix within the block:
@@
where $w$ and $u$ are two trainable parameters. The parameter $u$ is a bonus that rewards the model for encountering a token for the first time, specifically the current token. This helps the model pay more attention to the current token and circumvents any potential degradation of $w$. Another important parameter is $w$, which is a channel-wise time decay vector per head. Furthermore, we transform parameter $w$ by $w=\exp(-\exp(w))$. This transformation ensures that all values of ${w}$ are within the range $(0,1)$, ensuring that $\mathrm{diag}(w)$ represents a contraction matrix.
@@ 3.2.2 Data-dependent Recurrence
The key improvement over data-independent time mixing (Eq. 2) lies in the evolution of the time decay vector from a fixed parameter $w$ to a dynamic one $w_{t}$ that reacts to the input data. The dynamic nature of $w_{t}$ allows the model to adjust more nimbly to diverse input data, unbound by rigid, predefined structures. Equations are as follow:
@@
The LoRA mechanism utilizes vectors learned from data-independent time mixing and enhances them at a low cost with additional offsets determined by the incoming input. It should be noted that the computation of the new time-varying decay $w_{t}$ employs a token-shifted value as its input, not just the current token. As shown in Table 1, the VisualRWKV equipped with data-dependent recurrence exhibits a significant improvement in performance.
@@ 3.2.3 Sandwich Prompt
RWKV and other linear RNN models are prompt-sensitive. Therefore, designing specific prompts to assist VisualRWKV in acquiring information is crucial. For this purpose, we have specifically designed three types of prompting methods, as shown in Figure 2:
@@
The Sandwich Prompt demonstrates the best performance, as it provides a stronger conditioning for model generation.
@@ 3.2.4 Image Scanning
Since vanilla RWKV is designed to handle 1D sequential data with causal relationships, such as language sequences, while the visual sequences generated by the vision encoder are non-causal data, this work proposes a 2D scanning mechanism to enhance visual task performance. This research is dedicated to enhancing the performance of VisualRWKV by integrating the 2D scanning mechanism into the RWKV blocks. Specifically, we explore three variants of multimodal RWKV blocks as shown in Figure 4:
@@ 4.4.2 Ablation on Prompting Method
As shown in Table 3, of the three prompting approaches, our proposed sandwich prompt performs the best, with the conventional image-first prompt coming in second, and the image-last prompt being the least effective. The enhanced effectiveness of the sandwich prompt can be elucidated as such: by enabling the model to first review the instructions prior to engaging with the image, the sandwich prompt facilitates a more targeted extraction of information from the image, thereby strengthening the conditional aspects of the image information retrieval process. However, merely placing the instructions before the image is insufficient. We’ve observed that the image-last prompt is notably less effective. Placing the instructions merely ahead of the image is inadequate; we’ve found that the image-last prompt is notably less effective. This occurs because Linear RNN models tend to forget the instructional information following image processing, necessitating a repetition of the instructions to attain improved outcomes. Furthermore, our research indicates that the sandwich prompt is capable of efficiently mitigating information loss due to a decrease in image tokens, maintaining good results with only a small number of image tokens. Additional experimental outcomes and detailed analysis are available in Appendix E.
@@ 4.4.3 Ablation on Scanning Method
Table 4: Results for three scanning method.
@@ 4.5 Efficiency Analysis
As shown in Figure 1, we compared the inference speed and GPU memory consumption directly with LLaVA-1.5 of the same parameter size. VisualRWKV has a constant single token inference speed, while the inference speed of a single token in LLaVA-1.5 slows down as more tokens are generated. On the other hand, VisualRWKV has a constant GPU memory consumption, while the memory consumption of LLaVA-1.5 increases linearly. In practice, compared to LLaVA-1.5, VisualRWKV has a speed advantage of 3.98 times and can save 54% of the GPU memory when reaching an inference length of 24576 tokens.
@@ 5 Conclusions
Despite the encouraging results achieved by VisualRWKV, several limitations must be acknowledged. Firstly, due to the lack of data that follows such instructions and the limitation of context length, VisualRWKV is still unable to process multiple images. Secondly, despite VisualRWKV shows good performance on academic datasets, its capacity to address problems in certain areas like TextVQA could be constrained by the limitations in the recall ability of efficient Language models (Arora et al., 2023). This can be enhanced through further improvements to the model’s architecture. Lastly, in order to be consistent with LLaVA-1.5, this study did not explore the impact of the vision encoder and training data on VisualRWKV. In the future, we will attempt powerful visual encoder and use higher-quality training data to further enhance the performance of VisualRWKV.
@@
Although VisualRWKV significantly reduces the occurrence of hallucinations, it can still generate hallucinations and occasionally disseminate misinformation. Therefore, its use in critical fields such as the medical industry must be approached with extreme caution.

### v2

@@ Abstract
Visual Language Models (VLMs) have rapidly progressed with the recent success of large language models. However, there have been few attempts to incorporate efficient linear Recurrent Neural Networks (RNNs) architectures into VLMs. In this study, we introduce VisualRWKV, the first application of a linear RNN model to multimodal learning tasks, leveraging the pre-trained RWKV language model. We propose a data-dependent recurrence and sandwich prompts to enhance our modeling capabilities, along with a 2D image scanning mechanism to enrich the processing of visual sequences. Extensive experiments demonstrate that VisualRWKV achieves competitive performance compared to Transformer-based models like LLaVA-1.5 on various benchmarks. Compared to LLaVA-1.5, VisualRWKV has a speed advantage of 3.98 times and can save 54% of GPU memory when reaching an inference length of 24K tokens. To facilitate further research and analysis, we have made the checkpoints and the associated code publicly accessible at the following GitHub repository: https://github.com/howard-hou/VisualRWKV.
@@
Haowen Hou  and Peigen Zeng  and Fei Ma  and Fei Richard Yu Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ), Shenzhen, China {houhaowen, mafei, yufei}@gml.ac.cn
@@ 3.1 Preliminaries
As shown in Figure 3, trainable variable $\mu_{g}$, $\mu_{r}$, $\mu_{k}$, $\mu_{v}$ are used in a linear combination of $x_{t}$ and $x_{t-1}$, to achieve a simple mixing, which interpolate between the inputs of the current and previous time-steps. The combination of shifted previous step and current step was linear projected through projection matrix within the block:
@@
where $w$ and $u$ are two trainable parameters. The parameter $u$ serves as a term weight for the current token when the model encounters it for the first time. It enables the model to efficiently process the token by focusing more on important tokens and quickly filtering out unimportant ones. Another important parameter is $w$, which is a channel-wise time decay vector per head. Furthermore, we transform parameter $w$ by $w=\exp(-\exp(w))$. This transformation ensures that all values of ${w}$ are within the range $(0,1)$, ensuring that $\mathrm{diag}(w)$ represents a contraction matrix.
@@ 3.2.2 Data-dependent Recurrence
The key improvement over data-independent time mixing (Eq. 2) lies in the evolution of the time decay vector from a fixed parameter $w$ to a dynamic one $w_{t}$ that reacts to the input data $x_{t}$ at time step $t$. The dynamic nature of $w_{t}$ allows the model to adjust more nimbly to diverse input data, unbound by rigid, predefined structures. Equations are as follow:
@@
The LoRA mechanism utilizes vectors learned from data-independent time mixing and enhances them at a low cost with additional offsets modulated by the incoming input. It should be noted that the computation of the new time-varying decay $w_{t}$ employs a token-shifted value $\mathrm{ddlerp}_{d}(x_{t},x_{t-1})$ as its input, not just the current token $x_{t}$. As shown in Table 1, the VisualRWKV equipped with data-dependent recurrence exhibits a significant improvement in performance.
@@ 3.2.3 Sandwich Prompt
The motivation for designing the sandwich prompt is as follows: Unlike the attention mechanism in Transformers, RNN models such as RWKV, due to their sequential nature, cannot revisit historical information repeatedly. Instead, they must decide immediately whether to store a token or image token in memory upon encountering it. Therefore, carefully designing tailored prompts is essential for enhancing VisualRWKV’s ability to effectively acquire and utilize information. For this purpose, we have specifically designed three types of prompting methods, as shown in Figure 2:
@@
The sandwich prompt is designed to provide optimal conditions that assist the model in making these decisions more effectively. Specifically, the first prompt helps the model efficiently extract relevant information from the image, while the second prompt focuses on improving the model’s ability to answer questions.
For instance, the Image Last Prompt can cause the model to occasionally forget the question embedded in the prompt, while the Image First Prompt may result in the model processing the image without considering the question, hindering its ability to analyze the image contextually. In contrast, the sandwich prompt resolves these issues and achieves a synergistic effect, enabling the model to perform better than the sum of the individual prompts. The experimental results show that the Sandwich Prompt achieves the best performance, as presented in Table 3.
@@ 3.2.4 Image Scanning
The motivation for designing the image scanning techniques is as follows: Language is inherently unidirectional, while images are multidirectional by nature. As a result, unidirectional language models face inherent limitations when processing visual information. By implementing bidirectional or multidirectional image scanning strategies, these challenges can be effectively mitigated.
Vanilla RWKV is designed for 1D sequential data with causal relationships, such as language sequences. However, the visual sequences generated by vision encoders are non-causal. To bridge this gap, we propose a 2D scanning mechanism to improve VisualRWKV’s performance on visual tasks. This work integrates the 2D scanning mechanism into RWKV blocks, exploring three variants of multimodal RWKV blocks, which are illustrated in Figure 4:
@@
Our design alternates different scanning directions within layers, which does not introduce additional computational overhead and preserves the efficiency of the architecture. The experimental results (Table 4) have also verified the effectiveness and necessity of such scanning techniques in enhancing the model’s ability to handle and understand visual sequences, thereby improving the overall performance of VisualRWKV in visual language processing tasks.
@@ 4.3.1 Main Results
4.3.2 Gain Analysis on Different Benchmarks
VisualRWKV excels in academic benchmarks like VQA, GQA, and SQA, where both the questions and answers are short texts. The model faces no fundamental obstacles in handling such tasks, leading to significant performance improvements. As a result, VisualRWKV achieves results that are comparable to, and even surpass, the Transformer-based LLaVA-1.5 on these benchmarks.
Although VisualRWKV shows notable improvements on the TextVQA (TQA) benchmark, it still lags behind LLaVA-1.5 in this task (51.0 vs. 58.2). TextVQA requires recalling information from images, which is similar to the Multi-Query Associative Recall (MQAR) task Arora et al. (2023), which is often a limitation for RNN-like architectures. However, our latest work, VisualRWKV-HD/UHD Li and Hou (2024), has shown that higher resolution and better-quality image features can significantly alleviate these limitations.
@@ 4.4.2 Ablation on Prompting Method
As shown in Table 3, among the three prompting approaches, the sandwich prompt outperforms the others, followed by the image-first prompt, with the image-last prompt being the least effective. The effectiveness of the sandwich prompt is attributed to its ability to allow the model to review the instructions before engaging with the image, enabling a more targeted extraction of information and enhancing the conditional aspects of image information retrieval.
However, simply placing the instructions before the image is insufficient. The image-last prompt performs poorly because linear RNN models tend to forget the instructions after processing the image, making it necessary to repeat the instructions for better results. Additionally, our research shows that the sandwich prompt can effectively mitigate information loss even with a reduced number of image tokens, maintaining robust performance. Further experimental results and analyses can be found in Appendix E.
@@ 4.4.3 Ablation on Scanning Method
The image scanning techniques are applied during both training and inference, and it is essential to maintain train-test consistency. We have made simple attempts to rearrange the order of layers with different directions, but the performance was not robust. Specific layers have already been specialized to process image information from particular directions.
@@
Table 4: Results for three scanning methods.
@@ 4.5 Efficiency Analysis
As shown in Figure 1, we compared the inference speed and GPU memory consumption directly with LLaVA-1.5 of the same parameter size. VisualRWKV has a constant single token inference speed, while the inference speed of a single token in LLaVA-1.5 slows down as more tokens are generated. On the other hand, VisualRWKV has a constant GPU memory consumption, while the memory consumption of LLaVA-1.5 increases linearly. In practice, compared to LLaVA-1.5, VisualRWKV has a speed advantage of 3.98 times and can save 54% of the GPU memory when reaching an inference length of 24576 tokens. Since VisualRWKV retains a fixed state size throughout inference, GPU memory usage remains nearly constant, which is illustrated as a straight line in Figure 1(c).
@@ 5 Conclusions
Despite the encouraging results achieved by VisualRWKV, several limitations must be acknowledged. Firstly, due to the lack of data following such instructions and the limited context length, VisualRWKV is currently unable to process multiple images. Secondly, although VisualRWKV shows good performance on academic datasets, its ability to handle certain tasks, such as TextVQA, may be constrained by the limitations in the recall ability of efficient language models (Arora et al., 2023). These constraints can potentially be mitigated by further architectural improvements. Lastly, to maintain consistency with LLaVA-1.5, this study did not investigate the effects of the choice of vision encoder or the quality of training data on VisualRWKV. In the future, we aim to explore more advanced visual encoders and utilize higher-quality training data to further enhance its performance.
@@
Although VisualRWKV significantly reduces the occurrence of hallucinations, it can still generate hallucinations and occasionally disseminate misinformation. Therefore, its application in critical fields, such as the medical industry, should be approached with great caution.
@@ Acknowledgments
Supplementary Material for VisualRWKV: Exploring Recurrent Neural Networks for Visual Language Models